# TritonRL: Training LLMs to Think and Code Triton Without Cheating

### v1

@@ Appendix D Metrics
Appendix E GRPO with uniform reward assignment
@@
$\begin{aligned} \mathcal{J}_{\text{GRPO}}(\theta)&=\mathbb{E}_{q\sim P(Q),~\{o_{i}=(o_{i}^{\text{plan}},o_{i}^{\text{code}})\}_{i=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q)}\left[\frac{1}{G}\sum_{i=1}^{G}\mathcal{L}_{\text{GRPO}}(\theta,i)\right]\end{aligned}$
where $\pi_{\theta}$ and $\pi_{\theta_{old}}$ are the policy model and reference model, $q$ denotes a prompt given to the model, defining a task to implement in Triton, and $o_{i}$ represents $i$-th response generated by the model for $q$ in the group $G$. The GRPO losses $\mathcal{L}(\theta,i)$ is computed as
$\begin{aligned} \mathcal{L}_{\text{GRPO}}(\theta,i)&=\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}A_{i}\cdot\min\left\{\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})},\text{clip}\left(\frac{\pi_{\theta}(o_{i,t}|q,o_{i,<t})}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i,<t})},1-\epsilon,1+\epsilon\right)\right\},\end{aligned}$
where $A_{i}$ denotes the group-wise advantage uniformly applied to all tokens, computed as $A_{i}=r_{i}-\frac{1}{G}\sum_{j=1}^{G}r_{j}$, with $r_{i}$ being the reward for the $i$-th sample. In the uniform reward assignment, we define the reward as a weighted combination of correctness and speedup conditioned on validity, which can be expressed as
$r_{i}=\texttt{syntax}_{i}\cdot\texttt{func}_{i}\cdot(\beta\cdot\texttt{correct}_{i}+(1-\beta)\cdot\texttt{speedup}_{i}),$ (8)
@@
Appendix F Data curation and examples
F.1 Data Mixing Subset Creation
@@
Data Mixing for Reinforcement Learning.
In our task, we have a training dataset $\mathcal{D}_{train}$ that is made of examples from KernelBook. We have an evaluation dataset $\mathcal{D}_{test}$ from KernelBench. We regard the problem of selecting the right examples to post-train our LLM model as a data mixing optimization problem, and we formulate the problem as follows.
Given a training dataset $\mathcal{D}_{train}$ and a test dataset $\mathcal{D}_{test}$, either dataset can be divided into $m$ subsets. we suppose the probability of selecting $m$ training data subsets forms a probability simplex $\mathbf{p}\in\Delta$, which represents the probability of drawing different subsets in the training set. We define the reward function $R_{test}^{i}(\bm{p})=\frac{1}{|\mathcal{D}_{test}^{i}|}\sum_{q\in\mathcal{D}_{test}^{i}}\max_{o\in\pi^{\bm{p}}(q)}r\big(q,o\big)$, where $\pi^{\bm{p}}$ is the policy that is trained under the data mixture $\bm{p}$. In this problem, we use a static mixture $\bm{p}$ for all post-training steps. Following the proposal by Chen et al. (2024), we can express a data mixture optimization problem as the following optimization problem:
$\begin{aligned} \underset{{\bm{p}\in\Delta}}{\text{maximize}}\;\sum_{j=1}^{m}R_{test}^{j}(\bm{p})\;\text{s.t.}\;R_{test}^{j}(\bm{p)}=\zeta(S,\bm{p}),\text{where}\;S\in\mathbb{R}^{m\times m}.\end{aligned}$
We define $\mathcal{S}$ as the reward interaction matrix, where ${\mathcal{S}}_{ij}$ captures the effect of post-training on the $i$-th subset in $\mathcal{D}_{train}$ and then using the learned policy $\pi$ to generate corresponding kernels for the $j$-th $\mathcal{D}_{test}$. We use a general function $\zeta$ to leverage this information to predict the reward obtained on the $j$-th subset in $\mathcal{D}_{test}$.
We make two crucial changes to prior data mixing work. First, unlike the typical data mixing work that assumes $\zeta$ can be parametrized as a linear function, we do not make specific assumptions (Chen et al., 2023; Xie et al., 2023; Fan et al., 2023). Second, we do not create random subsets of our data. Instead, we note that KernelBench (Ouyang et al., 2025) has three complexity levels. We use an LLM labeler to create difficulty labels for each data point in $\mathcal{D}_{train}$. Since we focus on difficulty level 1 and 2, we create two subsets for training and test: $\mathcal{D}_{train}=\{\mathcal{D}_{train}^{L1},\mathcal{D}_{train}^{L2}\}$, $\mathcal{D}_{test}=\{\mathcal{D}_{test}^{L1},\mathcal{D}_{test}^{L2}\}$. Since the problem is under-constrained without a known parametric form of $\zeta$, we simply evaluate three candidate initializations, $\bm{p}\in{[1,0],[0,1],[0.5,0.5]}$, and choose the best-performing mixture rather than fully modeling $S$ and $\zeta$ or solving for $p$ exhaustively.
F.2 Example of invalid Triton codes in KernelBook

### v2

@@ Appendix C Metrics
To measure correctness, we execute the generated Triton code for five randomly generated inputs and compare the outputs against those from the reference PyTorch code. If all outputs match, we consider the generated code correct. To measure speedup, we executes a reference code and generated Triton code on the same hardware setup (NVIDIA L40S) for 10 times for a randomly generated input and compute the ratio of their mean execution times.
Appendix D GRPO with uniform reward assignment
@@
$\begin{aligned} \mathcal{J}_{\text{\tiny GRPO}}(\theta)&=\mathbb{E}_{q_{i}\sim\mathcal{D}_{RL},~\{o_{i,j}\}_{j=1}^{G}\sim\pi_{\theta_{old}}(\cdot|q_{i})}\left[\mathcal{L}_{\text{\tiny GRPO}}(\theta,i)\right]\end{aligned}$
where $\pi_{\theta}$ and $\pi_{\theta_{old}}$ are the policy model and reference model, $q_{i}$ denotes a prompt given to the model, defining a task to implement in Triton, and $o_{i,j}$ represents $i$-th response generated by the model for $q_{i}$ in the group $G$. The GRPO losses $\mathcal{L}(\theta,i)$ is computed as
$\begin{aligned} \mathcal{L}_{\text{\tiny GRPO}}(\theta,i)&=\frac{1}{G}\sum_{j=1}^{G}\frac{1}{|o_{i,j}|}\sum_{t=1}^{|o_{i,j}|}A_{i,j}\cdot\min\left\{\frac{\pi_{\theta}(o_{i,j,t}|q,o_{i,j,<t})}{\pi_{\theta_{old}}(o_{i,j,t}|q,o_{i,j,<t})},\text{clip}\left(\frac{\pi_{\theta}(o_{i,j,t}|q,o_{i,j,<t})}{\pi_{\theta_{old}}(o_{i,j,t}|q,o_{i,j,<t})},1-\epsilon,1+\epsilon\right)\right\},\end{aligned}$
where $A_{i,j}$ denotes the group-wise advantage uniformly applied to all tokens, computed as $A_{i,j}=r_{i,j}-\frac{1}{G}\sum_{j=1}^{G}r_{i,j}$, with $r_{i,j}$ being the reward for the $i$-th sample. In the uniform reward assignment, we define the reward as a weighted combination of correctness and speedup conditioned on validity, which can be expressed as
$\begin{aligned} r_{i,j}&=\beta\cdot R^{\text{correct}}(g_{i},o_{i,j})+(1-\beta)\cdot R^{\text{speedup}}(g_{i},o_{i,j}),\end{aligned}$
@@
Appendix E Training Hyperparameters and Details
For generating reasoning and code traces for SFT, we use temperature=0.6, top_p=0.95 for DeepSeek-R1 (Guo et al., 2025) and temperature=1.0, top_p=1.0 for GPT-OSS 120B (OpenAI, 2025). To label difficulty, we further label tasks into three difficulty levels using Qwen3-235B-Instruct (Team, 2025) (temperature=0.7, top_p=0.8).
For SFT, we train for 2 epochs on DeepSeek-R1-generated data with a maximum sequence length of $12,288$ tokens and 3 epochs on GPT-OSS 120B-generated data with a maximum sequence length of $16,384$ tokens, all with a batch size of 16 and a learning rate of $1\times 10^{-5}$.
For RL, we train the model for 2 epochs using the VeRL framework (Sheng et al., 2025) with a batch size of 32, a learning rate of $1\times 10^{-6}$, a clip ratio of $\epsilon=0.2$ for the GRPO objectives, the maximum prompt length $2,048$, and the maximum response length $16,384$. We use 8 NVIDIA A100 80GB GPUs for both SFT and RL training.
Note that TritonRL (G), a model trained with SFT checkpoint distilled from GPT-OSS 120120B, was trained with 8 NVIDIA H200 GPUs due to resource constraints. Also, we use KernelBook (Paliskara and Saroufim, 2025) tasks where input generating functions are augmented using GPT-OSS 120120B for RL training.
Appendix F Additional Experiment Results
F.1 Pass@k Results
We evaluated inference scaling by varying the number of sampled attempts ($k=1,5,10$), as illustrated in Figure 4 and Figure 5. The correctness and runtime efficiency of TritonRL increases with more samples, whereas KernelLLM and Qwen3-8B show limited improvement, suggesting that TritonRL generates a more diverse set of codes and benefits from additional sampling during inference.
Figure 4: Pass@$k$ correctness and $\text{fast}_{1}$ for $k=1,5,10$ on KernelBench Level 1 tasks. The figures show how performance (correctness on the left, $\text{fast}_{1}$ on the right) of TritonRL and baseline models scales as the number of attempts increases.
Figure 5: Pass@$k$ correctness and $\text{fast}_{1}$ for $k=1,5,10$ on KernelBench Level 2 tasks. The figures show how performance (correctness on the left, $\text{fast}_{1}$ on the right) of TritonRL and baseline models scales as the number of attempts increases.
We also report pass@1 and pass@5 results for KernelBench Level 1 and Level 2 tasks in Table 5 and Table 6. These results further demonstrate the strong performance of TritonRL in generating valid, correct, and efficient Triton code across various pass@$k$ metrics.
Model pass@1 pass@5 pass@10
Qwen3 (8B) $70.0$ $65.0$ / $10.0$ $3.0$ / $2.0$ $96.0$ $96.0$ / $17.0$ $5.0$ / $4.0$ $99.0$ $99.0$ / $23.0$ $9.0$ / $5.0$
Qwen3 (14B) $80.0$ $73.0$ / $17.0$ $4.0$ / $2.0$ $99.0$ $96.0$ / $23.0$ $7.0$ / $6.0$ $100.0$ $99.0$ / $35.0$ $9.0$ / $7.0$
Qwen3 (32B) $81.0$ $73.0$ / $27.0$ $7.0$ / $5.0$ $99.0$ $96.0$ / $53.0$ $32.0$ / $11.0$ $100.0$ $100.0$ / $62.0$ $39.0$ / $16.0$
AutoTriton $65.0$ $57.0$ / $35.0$ $8.0$ / $4.0$ $86.0$ $80.0$ / $54.0$ $19.0$ / $7.0$ $97.0$ $92.0$ / $57.0$ $25.0$ / $10.0$
TritonRL (D) $88.0$ $84.0$ / $44.0$ $19.0$ / $8.0$ $100.0$ $100.0$ / $70.0$ $29.0$ / $12.0$ $100.0$ $100.0$ / $78.0$ $36.0$ / $13.0$
TritonRL (D) + IA $88.0$ $83.0$ / $37.0$ $12.0$ / $9.0$ $100.0$ $99.0$ / $69.0$ $28.0$ / $14.0$ $100.0$ $99.0$ / $78.0$ $34.0$ / $15.0$
   SFT w. DeepSeek-R1 (w/o RL) $75.0$ $70.0$ /$31.0$ $13.0$ / $7.0$ $99.0$ $98.0$ / $49.0$ $25.0$ /$10.0$ $100.0$ $100.0$ / $54.0$ $28.0$ / $12.0$
TritonRL (G) $69.0$ $64.0$ / $25.0$ $12.0$ / $9.0$ $100.0$ $98.0$ / $59.0$ $28.0$ / $15.0$ $100.0$ $100.0$ / $76.0$ $35.0$ /$17.0$
TritonRL (G) + IA $72.0$ $72.0$ / $29.0$ $13.0$ / $10.0$ $100.0$ $100.0$ / $77.0$ $36.0$ / $22.0$ $100.0$ $100.0$ / $88.0$ $41.0$ / $22.0$
   SFT w. GPT-OSS 120B (w/o RL) $43.0$ $43.0$ / $15.0$ $6.0$ / $5.0$ $94.0$ $92.0$ / $48.0$ $19.0$ / $15.0$ $100.0$ $100.0$ / $64.0$ $26.0$ / $18.0$
Claude-3.7 $79.0$ $74.0$ / $31.0$ $10.0$ / $5.0$ $100.0$ $100.0$ / $51.0$ $21.0$ / $11.0$ $100.0$ $100.0$ / $63.0$ $29.0$ / $16.0$
DeepSeek-R1 (685685B) $83.0$ $81.0$ / $32.0$ $13.0$ / $5.0$ $99.0$ $98.0$ / $56.0$ $25.0$ / $12.0$ $100.0$ $100.0$ / $69.0$ $28.0$ / $13.0$
GPT-oss (120120B) $79.0$ $78.0$ / $46.0$ $18.0$ / $10.0$ $100.0$ $100.0$ / $69.0$ $30.0$ / $18.0$ $100.0$ $100.0$ / $81.0$ $33.0$ / $22.0$
Model pass@1 pass@5 pass@10
Qwen3 (8B) $7.0$ $7.0$ / $0.0$ $0.0$ / $0.0$ $29.0$ $29.0$ / $0.0$ $0.0$ / $0.0$ $38.0$ $37.0$ / $1.0$ $0.0$ / $0.0$
Qwen3 (14B) $5.0$ $5.0$ / $0.0$ $0.0$ / $0.0$ $26.0$ $24.0$ / $0.0$ $0.0$ / $0.0$ $40.0$ $38.0$ / $1.0$ $0.0$ / $0.0$
Qwen3 (32B) $10.0$ $10.0$ / $1.0$ $1.0$ / $1.0$ $28.0$ $26.0$ / $2.0$ $2.0$ / $1.0$ $36.0$ $35.0$ / $2.0$ $2.0$ / $1.0$
KernelLLM $0.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $1.0$ $0.0$ / $0.0$ $0.0$ / $0.0$ $3.0$ $0.0$ / $0.0$ $0.0$ / $0.0$
AutoTriton $1.0$ $1.0$ / $0.0$ $0.0$ / $0.0$ $5.0$ $5.0$ / $0.0$ $0.0$ / $0.0$ $11.0$ $10.0$ / $1.0$ $0.0$ / $0.0$
TritonRL (D) $28.0$ $27.0$ / $2.0$ $1.0$ / $0.0$ $55.0$ $55.0$ / $16.0$ $7.0$ / $1.0$ $63.0$ $63.0$ / $23.0$ $12.0$ / $3.0$
TritonRL (D) + IA $24.0$ $24.0$ / $3.0$ $3.0$ / $2.0$ $42.0$ $41.0$ / $14.0$ $10.0$ / $5.0$ $52.0$ $51.0$ / $24.0$ $16.0$ / $7.0$
   SFT w. DeepSeek-R1 (w/o RL) $22.0$ $19.0$ / $4.0$ $4.0$ / $0.0$ $50.0$ $45.0$ / $10.0$ $9.0$ / $2.0$ $55.0$ $51.0$ / $13.0$ $10.0$ / $3.0$
TritonRL (G) $49.0$ $47.0$ / $4.0$ $3.0$ / $1.0$ $88.0$ $87.0$ / $16.0$ $11.0$ / $2.0$ $91.0$ $91.0$ / $23.0$ $15.0$ / $7.0$
TritonRL (G) + IA $44.0$ $44.0$ / $6.0$ $5.0$ / $2.0$ $83.0$ $82.0$ / $18.0$ $14.0$ / $6.0$ $93.0$ $93.0$ / $28.0$ $19.0$ / $10.0$
   SFT w. GPT-OSS 120B (w/o RL) $41.0$ $40.0$ / $4.0$ $4.0$ / $0.0$ $80.0$ $79.0$ / $10.0$ $7.0$ / $1.0$ $88.0$ $88.0$ / $13.0$ $9.0$ / $3.0$
Claude-3.7 $16.0$ $13.0$ / $2.0$ $0.0$ / $0.0$ $32.0$ $31.0$ / $12.0$ $3.0$ / $0.0$ $34.0$ $34.0$ / $14.0$ $4.0$ / $1.0$
DeepSeek-R1 $7.0$ $5.0$ / $1.0$ $0.0$ / $0.0$ $24.0$ $23.0$ / $7.0$ $4.0$ / $3.0$ $31.0$ $31.0$ / $14.0$ $9.0$ / $4.0$
GPT-oss (120120B) $9.0$ $9.0$ / $2.0$ $1.0$ / $1.0$ $30.0$ $29.0$ / $9.0$ $6.0$ / $2.0$ $39.0$ $39.0$ / $15.0$ $12.0$ / $4.0$
F.2 Robustness gains from diverse input augmentation (IA)
To demonstrate the effectiveness of our input augmentation (IA) strategy in enhancing model robustness to diverse input shapes, we conducted additional experiments on input-augmented KernelBench Level 1 and Level 2 tasks, where five distinct input shapes were used for each of the 100 tasks per level (500 total evaluations per level). The results, presented in Table 7, indicate IA improves correctness and speedup for small $k$ (pass@1 and pass@5), showing its role in enhancing robustness to input variations on initial attempts. As $k$ increases to 10, the performance gap narrows or even reverses, suggesting that with sufficient sampling models can produce high-quality outputs regardless of exposure to diverse input shapes during training. Nonetheless, IA remains beneficial for robustness, particularly in low-sample scenarios.
Table 7: Pass@k results (k = 1, 5, 10) on input-augmented KernelBench Level 1 and 2. Each of 100 tasks for each level was evaluated with five distinct input shapes (500 total for each level).
Level 1
Model pass@1 pass@5 pass@10
TritonRL (D) $88.0$ $85.0$ / $37.0$ $12.0$ / $6.0$ $100.0$ $100.0$ /$64.0$ $27.0$ / $10.0$ $100.0$ $100.0$ / $75.0$ $33.0$ / $12.0$
TritonRL (D) + IA $85.0$ $81.0$ / $39.0$ $14.0$ / $7.0$ $100.0$ $99.0$ / $65.0$ $27.0$ / $11.0$ $100.0$ $99.0$ / $72.0$ $32.0$ / $13.0$
TritonRL (G) $64.0$ $62.0$ / $30.0$ $12.0$ / $7.0$ $99.0$ $99.0$ / $67.0$ $26.0$ / $17.0$ $100.0$ $100.0$ / $80.0$ $31.0$ / $19.0$
TritonRL (G) + IA $77.0$ $77.0$ / $38.0$ $14.0$ / $8.0$ $99.0$ $99.0$ / $73.0$ $29.0$ / $18.0$ $100.0$ $100.0$ / $83.0$ $36.0$ / $20.0$
Level 2
Model pass@1 pass@5 pass@10
TritonRL (D) $31.0$ $29.0$ / $5.0$ $2.0$ / $1.0$ $56.0$ $54.0$ / $18.0$ $9.0$ / $3.0$ $66.0$ $65.0$ / $27.0$ $15.0$ / $5.0$
TritonRL (D) + IA $25.0$ $24.0$ / $5.0$ $3.0$ / $1.0$ $43.0$ $42.0$ / $13.0$ $8.0$ / $3.0$ $52.0$ $51.0$ / $20.0$ $13.0$ / $4.0$
TritonRL (G) $44.0$ $42.0$ / $5.0$ $3.0$ / $1.0$ $85.0$ $84.0$ / $15.0$ $11.0$ / $5.0$ $94.0$ $93.0$ / $23.0$ $14.0$ / $6.0$
TritonRL (G) + IA $56.0$ $54.0$ / $8.0$ $6.0$ / $3.0$ $88.0$ $86.0$ / $22.0$ $14.0$ / $7.0$ $94.0$ $93.0$ / $29.0$ $19.0$ / $10.0$
F.3 Ablation on base model choices
In this section, we investigate the impact of different base models for fine-tuning with our RL framework. We compare two base models: Qwen3-8B, which is our main base model used throughout this paper, and Seed-Coder-8B-Reasoning (Zhang et al., 2025), which is the base model of AutoTriton (Li et al., 2025b). As shown in Table 8, TritonRL fine-tuned from Seed-Coder-8B-Reasoning outperforms AutoTriton, which is also fine-tuned from the same base model, across all metrics on KernelBench Level 1 and Level 2 tasks. This demonstrates the effectiveness of our RL fine-tuning approach regardless of the base model. Furthermore, when comparing TritonRL fine-tuned from Qwen3-8B and from Seed-Coder-8B-Reasoning, we observe that the former achieves significantly higher correctness and speedup on both Level 1 and Level 2 tasks, indicating that Qwen3-8B is a more effective foundation for our RL fine-tuning.
Table 8: Comparison of TritonRL fine-tuned from different base models (Qwen3-8B and Seed-Coder-8B-Reasoning) on KernelBench Level 1 and Level 2. All metrics are reported as pass@10 (%). TritonRL (G) indicates our model initialized from the GPT-OSS 120B SFT checkpoint; + IA denotes input augmentation. ”Qwen3-8B + TritonRL (G) + IA” and ”Seed-Coder-8B-Reasoning + TritonRL (G) + IA” refer to models fine-tuned with our RL framework from the respective base models under identical settings. AutoTriton uses Seed-Coder-8B-Reasoning as its base model and can be directly compared to ”Seed-Coder-8B-Reasoning + TritonRL (G) + IA”.
Model #Params Level1 Level2
valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$ valid compiled / correct $\textbf{fast}_{1}$ / $\textbf{fast}_{2}$
Seed-Coder-8B-Reasoning (base) 88B $91.0$ $82.0$ / $19.0$ $7.0$ / $6.0$ $49.0$ $12.0$ / $1.0$ $1.0$ / $0.0$
AutoTriton 88B $97.0$ $92.0$ / $57.0$ $25.0$ / $10.0$ $11.0$ $10.0$ / $1.0$ $0.0$ / $0.0$
Seed-Coder-8B-Reasoning + TritonRL (G) + IA 88B 99.0 99.0 / 61.0 27.0 / 17.0 86.0 85.0 / 17.0 11.0 / 5.0
Qwen3-8B (base) 88B $99.0$ $99.0$ / $23.0$ $9.0$ / $5.0$ $38.0$ $37.0$ / $0.0$ $0.0$ / $0.0$
Qwen3-8B + TritonRL (G) + IA 88B $100.0$ $100.0$ / $88.0$ $41.0$ / $22.0$ $93.0$ $93.0$ / $28.0$ $19.0$ / $10.0$
F.4 Comparison of token-class reward assignments
In GRPO, the advantage for the ”code” component is computed relative to other codes sampled in the same group, but those codes may be paired with different prior reasoning traces (plans). Penalizing a correct implementation solely because it was conditioned on a weak plan is undesirable: it risks discouraging valid Triton implementations and degrading the base model’s already limited Triton skills. Thus, we assign correctness-based rewards for code tokens to avoid such unwarranted penalties and to preserve model’s capability to implement correct Triton code.
We compare our choice of reward assignment (defined in (2)),
$r_{i,j}^{\text{\tiny plan}}=R^{\text{speedup}}(g_{i},o_{i,j}),\quad r_{i,j}^{\text{\tiny code}}=R^{\text{correct}}(g_{i},o_{i,j}),$ (8)
against models trained using the intuitive speedup rewards for both plan and code tokens, i.e.,
$r_{i,j}^{\text{\tiny plan}}=R^{\text{speedup}}(g_{i},o_{i,j}),\quad r_{i,j}^{\text{\tiny code}}=R^{\text{speedup}}(g_{i},o_{i,j}),$ (9)
for the same $\alpha=0.1,$ starting from the same SFT base model distilled from DeepSeek-R1. We denote our default reward assignment ((2)) as speedup-correct and the latter ((9)) as speedup-speedup assignment in the Table 9.
As shown in Table 9, giving correctness feedback to code tokens yields better performance in both speedup and correctness than giving speedup-based feedback to all tokens. As explained above, correctness rewards for code tokens prevent accurate implementations conditioned on suboptimal plans from being unfairly penalized, ultimately helping the model learn to generate kernels that are more efficient.
Table 9: Comparison of token-class reward assignments (speedup-correct (ours) vs. speedup-speedup) on KernelBench Level 1 tasks. All metrics are reported as pass@10 (%).
reward assignment type valid compiled / correct $\textbf{fast}_{\textbf{1}}$ / $\textbf{fast}_{\textbf{2}}$
speedup-correct (ours) $100.0$ $100.0$ / $78.0$ $36.0$ / $13.0$
speedup-speedup $100.0$ $100.0$ / $78.0$ $33.0$ / $13.0$
Appendix G Data curation and examples
G.1 Data Mixing Subset Creation
@@
G.2 Example of invalid Triton codes in KernelBook